7. Gravetter/Wallnau/Forzano, Essentials Chapter 12 End-of-chapter question 22 Aa Aa 을 The following data have the same sample means and variances that appeared in the preceding question but the sample size is increased to n = 10. n 10 n=10 n 10 T-10 T=50 T=60 s2-9.00 s2 = 10.00 s2-11.00 SS-81 SS = 90 SS = 99 G=120 Σχε = 890 Predict how the increase in sample size should affect the F-ratio for these data. Use an ANOVA to check your prediction Larger samples should the F-ratio. Source df MS Between treatments Within treatments Total Predict how the increase in sample size should affect the value of η 2 for these data compared to the η2 in problem 10. Calculate n2 to check your prediction. Increasing the sample size should effect on η 2,
Answer #1

large samples should increase the F ratio

Group ni i S2i ni*(Xi-Xgrand)2 (ni-1)*S2i
1 10 1 9.000 90.000 81.00
2 10 5 10.000 10.000 90.00
3 10 6 11.000 40.000 99.00
grand mean= 4.0000 140.000 270.00
SSTr SSE
Source SS df MS F
between 140.00 2.000 70.0000 7.00
within 270.00 27.000 10.0000
total 410.00 29.0000

increasing the sample size should decrease the effect on \eta2

\eta2 =SSR/SST =140/410=0.3415